# Distributed network of smartphone sensors: a new tool for scientific field measurements

### v2

@@ II.1.3 Accelerometer &\&& gyroscope calibration
We conclude that in practice, the gyroscope is reliable only to measure pure rotational motions around the gyroscope sensor, and cannot be used in general in combination with the accelerometer to decompose arbitrary superposition of translational and rotational motions. However, in some limit cases, in particular when the center of rotation is known, we expect the gyroscope to be reliable. One may refer to Fig. 4(f) to evaluate the systematic error as a function of both the angular frequency and the distance $r$ between the sensor and the axis of rotation.
II.2 Smartphone fleet remote control
II.2.1 Remote control of a single smartphone
@@
Numerous smartphone applications exist to record the phone sensor data. We first use the application Phyphox Staacks et al. (2018) developed at the university of Aachen, which provides a user-friendly interface to acquire data from several sensors in parallel. Phyphox provides in particular a standard URL communication protocol to send instructions (START, STOP, CLEAR) from a distant host. The sensor data can also be downloaded remotely using the Phyphox URL protocol (SAVE). We parametrize Phyphox with a custom module that acquires simultaneously the accelerometer, gyroscope, magnetometer and GPS data at their respective maximum sampling frequencies. As described below in this section, we managed to use Phyphox simultaneously on a fleet of 60 smartphones to run acquisitions and gather data remotely. However, we have faced several limitations inherent to the Phyphox design. First, Phyphox does not continuously save the sensor data on the smartphone internal storage, such that the recordings accumulate in a buffer memory, limiting the duration of continuous recordings to a size of about 10MB without interruption. Second, Phyphox requires an initial physical access to the phone to activate the distant access. Third, no time synchronization protocol with a distant host is implemented. To circumvent these three limitations, we developed an Android smartphone application called Gobannos Casoli et al. (2025). This application allows remote access to the smartphone sensors, continuous saving of sensor data to the smartphone memory, and synchronization of the phone’s clock with a remote computer with a better accuracy.
II.2.2 Network configuration & remote control of the fleet
To operate the fleet remotely, we connect all smartphones to the same Wi-Fi network. We use a Wi-Fi broadcast system that supports at least 70 parallel connections. A laptop connected to this local network, namely the controller, is used as a DHCP server for Wi-Fi broadcast system. This controller is also used to send remote instructions to the smartphone fleet. The DHCP server attributes to each smartphone a static IP address 192.168.X.1YY, where YY is the smartphone identification number, and X is the local network identifier. The correspondence between the MAC address and the IP is saved in a text file, and used to write the configuration file of the DHCP server. The local network is disconnected from Internet to avoid pop-up messages. We notice that for each smartphone, the network parameters need to be set to static IP address to avoid Wi-Fi disconnections, when connected to a local network without internet connection. Eventually, we can identify each smartphone with its static IP address.
All the smartphones run on the same version of the Android operating system. In a early version, we use Android Debug Bridge (ADB) to communicate command line instructions to the smartphones from the controller. However, ADB protocol has to be authorized on the phone, and the procedure depend on the smartphone type and the constructor. We first unlock the developer mode manually on each smartphone. To authorize ADB USB connections, Xiaomi required a Mi account associated with a personal SIM card and an email address. Each series of 3 smartphones required a different combination of SIM card and email address to be unlocked, but up to 6 smartphones could be unlocked for each email address using two different SIM cards. We then used 22 different SIM cards and 11 generic email addresses to unlock the 66 smartphones. Once connected to a Mi account, we allow the ADB USB connections, and the SIM card can then be removed. This procedure is a priori specific to Redmi, as other constructors do not necessarily require personal information to authorize ADB USB connections.
Prior to each day of experiment, each smartphone was connected to the controller using an USB cable, and an ADB connection was set automatically using a script adb_usb.sh. The script uses a second table to map the ADB USB unique smartphone identifier to the right IP address. The ADB links are stable for at least a day, as long as the smartphone is connected continuously to the Wi-Fi. The individual ADB links between the controller and the smartphones are used to start Phyphox remotely, check the phone state (battery level, temperature, activity), and exchange files between the controller and the smartphones. Once the Phyphox communication protocol is enabled, we use URL instructions to start, stop, and save the data remotely. We conclude that ADB is a versatile toolbox for controlling the smartphone state, however, it can become time-consuming. Our later implementation of Gobannos have made the ADB link mostly obsolete for our purpose.
@@ II.2.2 Network configuration & remote control of the fleet
II.2.3 Time synchronization
The time synchronization of a distributed network such as a smartphone fleet is challenging. For synchronizing clocks in a distributed network, the Network Time Protocol (NTP) was introduced and normalised in 1985 and is used to synchronize computer clocks over Internet Mills (1991). For better performance, the Precision Time Protocol (PTP) Eidson, Fischer, and White (2002) was later introduced. Both protocols use information exchange between a server and a client, to estimate the time delay between their internal clocks. The delay introduced by the communication time between the two machines is deduced, by assuming that the communication time is symmetric, and does not depend on time. The typical precision of a NTP protocol on a local network is of the order of 1 ms, while a PTP protocol can achieve $\mu$s precision under optimal conditions.
To test the clock synchronization using standard protocol, we implemented manually a network time protocol (NTP) and performed $10^{5}$ time requests between a server and 3 test smartphones using ADB time requests. Due to standard delays in Wi-Fi communication protocol, the typical duration of a time request is 100 milliseconds. Overall, the error of the NTP was found to be of few ms for a optimal Wi-Fi communication (direct line, less than 3 meter distance), but rises up to 100 ms for larger Wi-Fi distance. In optimal conditions, we used these time requests to estimate the properties of the phone clock. We found that the typical delay between two phones is of the order of 1 second, a much larger delay than the precision of the GPS clock. However, the time difference between two phone clocks is stable over time: we identify less than 1 ms of time drift between two smartphone clocks over 10 hours. A time synchronization every day was then found to be sufficient. In practice, a mechanical time synchronization gives a typical error of few ms between the smartphone clocks, and was found to be sufficient for most applications.
@@ II.2.3 Time synchronization
II.2.4 PhoneFleet application
@@
To facilitate the remote control of the smartphone fleet, we developed a Python application called PhoneFleet available on github.com Casoli et al. (2024), which implements the functionnalities described above. It includes in particular the DHCP server configuration, the creation of ADB links, the start of Phyphox remotely, the control of Phyphox or Gobannos acquisitions through the URL protocol (run, stop clear and save). To optimize time delays in the execution of a command on multiple phones, PhoneFleet uses asynchronous Python functions, as well as multi-threading to run tasks in parallel, on subparts of the smartphone fleet. PhoneFleet currently uses a graphical interface based on PyQt5 with buttons and tabs to organize the functions in themes, and run the commands on the selected phones. The application format may evolve in a near future to improve portability and maintenance capacity.
III Applications
@@
Here we present two applications of the smartphone fleet as time synchronized IMU sensors. The first section is devoted to the oscillation of a smartphone chain immersed in a turbulent flow. Each phone is then used both as the object of interest and the measuring device. The second section is devoted to the use of smartphones as local wave buoys, when placed on sea ice covering water. Recording the acceleration, angular velocity and earth magnetic field, we manage to measure the wave amplitude at different locations.
III.1 Pendulum chain in turbulence
@@
The environmental artist Ned Kahn’s artworks “Kinetic Façade” are building facades covered by thousands of aluminium pendulum plates that oscillate harmoniously in the wind, creating regular patterns of ripples that resemble a fluttering flag Shelley and Zhang (2011) or sea waves Drazin (2002); Perrard et al. (2019). To gain in-depth understanding of such phenomena, recent experimental investigations by Zhang and Perrard Zhang and Perrard (2024) were conducted on oscillation measurement of a 1 meter chain of pendulum plates confronting a turbulent flow by camera imaging. They have evidenced wave-like advective patterns of pendulums oscillations, similar to Ned Kahn’s artworks. By spectral analyses in spatiotemporal Fourier space of the pendulum oscillations, it has been shown that these moving patterns could emerge in turbulent flow as a result of two distinct mechanisms. One can be described as a resonant response of each pendulum near the pendulum’s natural frequency of oscillation and the other one attributes to the direct response of the pendulums to the turbulent fluctuations. The maximum response is reached at the intersection between the two dispersion relations.
@@
To upscale the observable pendulum plate dynamics to a larger flow dimension comparable to that of a real facade, we used smartphones as rigid plate pendulums to measure spatiotemporal pressure fluctuations. We constructed a ten-meter-long chain of 60 uniformly spaced smartphones, each one hinged to a common rod along its top edge, allowing free oscillation around the $x$-axis. The built-in sensors, including accelerometers and gyroscopes, enable the determination of the instantaneous tilt angle of each smartphone. This setup provides highly temporally resolved measurements of oscillations and is space-efficient, especially where imaging techniques using cameras reach their limits.
@@
Fig. 6 shows the spatiotemporal chart of the tangential accelerations $a_{z}$ measured by the chain of smartphones. From left to right of the chart, we observe wavy patterns propagating across the chain with an approximately constant phase speed. We also observe interference patterns with reflected waves that propagate upstream. Two ten-second instantaneous time signals of tangential and radial accelerations are shown in Fig. 6 (inset). We see that the smartphone orientation is periodic at well defined frequencies through both signals. The radial acceleration (black curve) is seen to oscillate at a frequency twofold larger than that of the tangential acceleration (red curve). This frequency doubling originates from the contribution of the term $\ell_{S}\dot{\theta}^{2}$ where $\ell_{S}$ is the $y$ distance from the accelerometer to the rod and $\theta(t)$ the time dependent orientation of the smartphone.
Figure 6: a) Image of the wind tunnel measurement setup. The chain of smartphones is placed in the symmetry plane of the measuring section, in suspension with metal trestles. b) Spatiotemporal chart of the tangential acceleration $a_{z}$ of the chain of telephones, for a wind speed $U=9.3$ m/s. Inset: temporal signal of the tangential $a_{z}$ and radial acceleration $a_{y}$ of the smartphone #$\#50$ located at around $x=8$ m. Both the signal duration and smartphone position are indicated by the white dashed line.
III.2 Wave buoys

### v5

@@ 2.1.3 Accelerometer &\& gyroscope calibration
We conclude that in practice, the gyroscope of the redmi 10A is reliable only to measure pure rotational motions around the gyroscope sensor, and cannot be used in general in combination with the accelerometer to decompose arbitrary superposition of translational and rotational motions. However, in some limit cases, in particular when the center of rotation is known, we expect the gyroscope to be reliable. One may refer to Fig. 9(f) to evaluate the systematic error as a function of both the angular frequency and the distance $r$ between the sensor and the axis of rotation.
We have also conducted tests on a Fairphone 4, the results are shown in A. We found that the gyroscope was accurate up to the maximum tested angular frequency of 12rad/s. We deduce that the limitation we observed on the gyroscope sensor of the Redmi 10A are phone dependent, and each smartphone model must be tested before use, to benchmark the accuracy of their sensors in various configurations.
2.2 Smartphone fleet remote control
2.2.1 Remote control of a single smartphone
@@
Numerous smartphone applications exist to record the phone sensor data. We first use the application Phyphox [staacks2018advanced] developed at the university of Aachen, which provides a user-friendly interface to acquire data from several sensors in parallel. Phyphox provides in particular a standard URL communication protocol to send instructions (START, STOP, CLEAR) from a distant host. The sensor data can also be downloaded remotely using the Phyphox URL protocol (SAVE). We parametrize Phyphox with a custom module that acquires simultaneously the accelerometer, gyroscope, magnetometer and GPS data at their respective maximum sampling frequencies. As described below in this section, we managed to use Phyphox simultaneously on a fleet of 60 smartphones to remotely run acquisitions and gather data. However, we have faced several limitations inherent to Phyphox design. First, Phyphox does not continuously save the sensor data on the smartphone internal storage, such that the recordings accumulate in a buffer memory, limiting the duration of continuous recordings to a size of about 10MB without interruption. Second, Phyphox requires an initial physical access to the phone to activate the distant access. Third, no time synchronization protocol with a distant host is implemented. To circumvent these three limitations, we developed an Android smartphone application called Gobannos [Gobannos_2025]. This application allows remote access to the smartphone sensors, continuous saving of sensor data to the smartphone memory, and synchronization of the phone’s clock with a remote computer with a better accuracy.
2.2.2 Network configuration & remote control of the fleet
@@
To operate the fleet remotely, we connect all smartphones to the same Wi-Fi network. We use a Wi-Fi broadcast system that supports at least 70 parallel connections. A laptop connected to this local network, namely the controller, is used as a DHCP server for Wi-Fi broadcast system. This controller is also used to send remote instructions to the smartphone fleet. The DHCP server attributes to each smartphone a static IP address 192.168.X.1YY, where YY is the smartphone identification number, and X is the local network identifier. The correspondence between the MAC address and the IP is saved in a text file, and used to write the configuration file of the DHCP server. The local network is disconnected from the Internet to avoid unwanted notifications from the phone OEM applications. We notice that for each smartphone, the network parameters need to be set to static IP address to avoid Wi-Fi disconnections, when connected to a local network without internet connection. Eventually, we can identify each smartphone with its static IP address.
The scalability of the Wi-Fi broadcast system is mainly limited by network-level constraints. In practice, the maximum fleet size is set by the number of stable simultaneous Wi-Fi connections supported by the access point, as imposed by the manufacturer through hardware capabilities.
All the smartphones run on the same version of the Android operating system. In a early version, we use Android Debug Bridge (ADB) to communicate command line instructions to the smartphones from the controller. However, ADB has to be authorized on the phone, and the procedure depends on the smartphone type and the constructor. We first unlock the developer mode manually on each smartphone. To authorize ADB USB connections, Xiaomi required a Mi account associated with a personal SIM card and an email address. Each series of 3 smartphones required a different combination of SIM card and email address to be unlocked, but up to 6 smartphones could be unlocked for each email address using two different SIM cards. We then used 22 different SIM cards and 11 generic email addresses to unlock the 66 smartphones. Once connected to a Mi account, we allow the ADB USB connections, and the SIM card can then be removed. This procedure is a priori specific to Redmi, as other constructors do not necessarily require personal information to authorize ADB USB connections.
Prior to each day of experiment, each smartphone was connected to the controller using an USB cable, and an ADB connection was set automatically using the script adb_usb.sh. The script uses a second table to map the ADB USB identifier, unique to a smartphone, to the right IP address. The ADB links are stable for at least a day, as long as the smartphone is connected continuously to the Wi-Fi. The individual ADB links between the controller and the smartphones are used to start Phyphox remotely, check the phone state (battery level, temperature, activity), and exchange files between the controller and the smartphones. Once the Phyphox communication protocol is enabled, we use URL instructions to start, stop, and save the data remotely. We conclude that ADB is a versatile toolbox for controlling the smartphone state, however, it can become time-consuming. Our later implementation of Gobannos have made the ADB link mostly obsolete for our purpose.
2.2.3 Time synchronization
@@
To test the clock synchronization using standard protocol, we implemented manually a network time protocol (NTP) and performed $10^{5}$ time requests between a server and 3 test smartphones using ADB time requests. Due to standard delays in Wi-Fi communication protocol, the typical duration of a time request is 100 milliseconds. Overall, the error of the NTP was found to be of few ms for a optimal Wi-Fi communication (direct line, less than 3 meter distance), but rises up to 100 ms for larger distances to the Wi-Fi access point. In optimal conditions, we used these time requests to estimate the properties of the phone clock. We found that the typical delay between two phones is of the order of 1 second, a much larger delay than the precision of the GPS clock. However, the time difference between two phone clocks is stable over time: we identify less than 10 ms of time drift between two smartphone clocks over 10 hours. A time synchronization every day was then found to be sufficient. In practice, a mechanical time synchronization gives a typical error of few ms between the smartphone clocks, and was found to be sufficient for most applications.
[…]
2.2.4 Error estimate and fleet reliability
[…]
In practical conditions and in particular in outdoor environments, the most common source of error is the reliability of the Wifi network. Starting the acquisitions indoor before deployment allows us to achieve a 100% reliability. The ability of the Gobannos application to store the data locally at regular times also guarantees a 100% reliability on data storage. Starting, stopping and retrieving data from the phone in outdoor condition introduces additional sources of error, and we practically achieve a 95% success rate, the remaining phone requiring manual adjustment (restart Gobannos, Wi-Fi connection troubleshoot). The main error encountered originate from unstable Wi-Fi connection. The smartphone fleet has been used at a temperature ranging from -25 C to 30 C, and shows equal reliability in this range. Note that for negative temperature, each phone was embedded in a 3D printed box, that insulates thermally the phones, reaching typically a temperature $15^{\circ}$C higher than the external temperature.
accelerometer gyroscope magnetometer
Resolution 0.0012 $\mathrm{m/s^{2}}$ 0.0011 $\mathrm{rad/s}$ 0.15 $\mathrm{\mu T}$
Rate 403.23 $\pm$ 4.35 Hz 403.23 $\pm$ 4.35 Hz 50.0 $\pm$ 1.31 E-5 Hz
Average 9.82 $\pm$ 0.02 $\mathrm{m/s^{2}}$ 0.02 $\pm$ 0.02 $\mathrm{rad/s}$ 104.76 $\pm$ 72.32 $\mathrm{\mu T}$
Time sync UDP protocol: 60 $\mathrm{\mu s}$. Mechanical protocol: 59.5 $\pm$ 8 $\mathrm{\mu s}$
Table 1: Characteristics of embedded IMU sensors for Redmi 10 smartphone model 22011119UY. Values are obtained from static test of 50 smartphones. *: data obtained from Phyphox database https://phyphox.org/sensordb/
2.2.5 PhoneFleet application
@@
To facilitate the remote control of the smartphone fleet, we developed a Python application called PhoneFleet available on github.com [Phonefleet_2024], which implements the functionnalities described above. It includes in particular the DHCP server configuration, the creation of ADB links, the start of Phyphox remotely, the control of Phyphox or Gobannos acquisitions through the URL protocol (run, stop clear and save). To optimize time delays in the execution of a command on multiple phones, PhoneFleet uses asynchronous Python functions, as well as multi-threading to run tasks in parallel, on subparts of the smartphone fleet. PhoneFleet currently uses a graphical interface based on PyQt5 with buttons and tabs to organize the functions in themes, and run the commands on the selected phones. The application format may evolve in the near future to improve portability and maintenance capacity.
3 Applications
@@
Here we present two applications of the smartphone fleet as time synchronized IMU sensors. The first section is devoted to the oscillation of a smartphone chain immersed in a turbulent flow. Each phone is then used both as the object of interest and the measuring device. The second section is devoted to the use of smartphones as local wave buoys, placed on floating sea ice. By recording the acceleration, angular velocity and earth magnetic field, we manage to measure the wave amplitude at different locations.
3.1 Pendulum chain in turbulence
@@
The environmental artist Ned Kahn’s artworks Kinetic Facade are building facades covered by thousands of aluminium pendulum plates that oscillate harmoniously in the wind, creating regular patterns of ripples that resemble a fluttering flag [shelley2011flapping] or sea waves [drazin2002introduction, perrard2019turbulent]. To gain in-depth understanding of such phenomena, recent experimental investigations by Zhang and Perrard [zhang2024large] were conducted on oscillation measurement of a 1 meter chain of pendulum plates confronting a turbulent flow by camera imaging. They have evidenced wave-like advective patterns of pendulums oscillations, similar to Ned Kahn’s artworks. By spectral analyses in spatiotemporal Fourier space of the pendulum oscillations, it has been shown that these moving patterns could emerge in a turbulent flow as a result of two distinct mechanisms. One can be described as a resonant response of each pendulum near the pendulum’s natural frequency of oscillation, and the other one attributed to the direct response of the pendulums to the turbulent fluctuations. The maximum response is reached at the intersection between the two dispersion relations.
@@
To upscale the observable pendulum plate dynamics to a larger flow dimension comparable to that of a real facade, we used smartphones as rigid plate pendulums to measure spatiotemporal pressure fluctuations. We constructed a ten-meter-long chain of 60 uniformly spaced smartphones, each one hinged to a common rod along its top edge, allowing free oscillation around the $x$-axis. The built-in sensors, including accelerometers and gyroscopes, enable the determination of the instantaneous acceleration of each smartphone. This setup provides highly temporally resolved measurements of oscillations and is space-efficient, especially where imaging techniques using cameras reach their limits.
Figure 5: a) Image of the wind tunnel measurement setup. The chain of smartphones is placed in the symmetry plane of the measuring section, in suspension with metal trestles. b) Technical drawing of the rod-mount suspension system and its assembly. $\ell_{S}=50.6$ mm and $\ell_{G}=111.8$ $m$ indicate the rod to the accelerometer distance and the rod to the center of gravity distance. c) Spatiotemporal chart of the tangential acceleration $a_{z}$ of the chain of telephones, for a wind speed $U=9.3$ m/s. Inset: temporal signal of the tangential $a_{z}$ (blue) and radial (green) acceleration $a_{y}$ of the smartphone #50\#50 located at around $x=8$ m. Both the signal duration and smartphone position are indicated by the white dashed line.
@@
In the following, we aim to characterize the spatio-temporal dynamics of the wavy patterns along the smartphone chain, based on accelerometer data. Fig. 5b shows the assembly of one smartphone as a pendulum oscillating around the $x$ axis. The accelerometer is located at a distance $\ell_{S}=50.6$ mm from the rod center axis. Assuming that the smartphone oscillates only around the $x$ axis, the angular motion of a single smartphone in a non Galilean frame of reference induces a tangential and a radial accelerations $a_{z}$ and $a_{y}$ given by [dauphin2018physical]:
$a_{z}$ $=$ $-g\sin\theta-\ell_{S}\ddot{\theta},$ (1)
$a_{y}$ $=$ $g\cos\theta+\ell_{S}\dot{\theta}^{2},$ (2)
The tangential component $a_{z}$ corresponds to the acceleration measured by the sensor along the local $z$ axis. For a simple pendulum, both contributions in Eq.1 oscillate at the natural frequency of the pendulum $\omega_{0}=9.37$ rad/s, which we determined empirically from free oscillation measurements. However, the radial component $a_{y}$ is dominated by a second-harmonic response at $2\omega_{0}$ (arising from the $\cos\theta$ and $\dot{\theta}^{2}$ terms in Eq. 2). Although $a_{z}$ and $a_{y}$ are nonlinear functions of the oscillating angle $\theta(t)$, their form depends only on the local motion and not on the global spatial pattern of the smartphone chain. Consequently, in the following we characterize the spatiotemporal chain dynamics directly from the raw accelerometer signals.
@@
Fig. 5c shows the spatiotemporal chart of the tangential accelerations $a_{z}$ measured by the chain of smartphones. From left to right of the chart, we observe wavy patterns propagating across the chain with an approximately constant phase speed. We also observe interference patterns with reflected waves that propagate upstream.
Figure 6: Spatiotemporal spectrum of the normalized tangential acceleration $\log_{10}|\hat{a}_{z}|(f/f_{0},kL)$ of the telephone chain for three increasing wind speeds (a): $U=1.92$ m/s, (b): $U=4$ m/s, (c): $U=6.58$ m/s. Frequency and wavenumber normalized by the natural frequency of oscillation $f_{0}$ and the smartphone spacing distance $L$. Local maxima of $|\hat{a}_{z}|$ on the dispersion relation IIII for the low frequency range $f<f_{0}$ are indicated by green circles. Purple dashed lines correspond to the slopes given by the mean wind speed $U=2\pi f/k$. (d) Convection velocity $U_{c}$ measured by linear slope fit on the dispersion relation IIII as a function of the wind speed.
We now characterize the wave dynamics of the smartphone chain in the frequency and wavenumber domain. We perform the two-dimensional discrete Fourier transform in space and in time to convert the tangential acceleration $a_{z}(x,t)$ in physical space into $|\hat{a}_{z}(f,k)|$ in spectral space. For each discrete frequency $f_{i}$, the spectral amplitude $|\hat{a}_{z}(f_{i},k)|$ is normalized by its maximum over all wavenumbers, ensuring values between 0 and 1. Fig. 6(a-c) shows the frequency-wavenumber spectrum of the magnitude of the tangential acceleration $|\hat{a}_{z}|$ for three increasing wind speeds. We observe two distinct dispersion relations governing the oscillation dynamics, consistent with our earlier observations on Ned Kahn’s Kinetic Facade artworks [zhang2024large]. Branch I corresponds to oscillations of the smartphone chain around the natural frequency $f/f_{0}=1$, marginally dependent on wavenumber $k$. We attribute this dispersion relation to a resonant response mechanism of a single smartphone to the wind’s turbulent fluctuations in the vicinity of the natural frequency $f_{0}$. Branch II follows a linear dispersion relation through the origin ($f/f_{0}=0$, $kL=0$) and describes the telephone oscillations at low frequencies $f<f_{0}$. These motions can be understood as a passive response to coherent turbulent structures convected by the wind at a characteristic velocity $U_{c}$, given by the slope of the branch. For comparison, the wind free-stream velocity $U$ is plotted as dashed purple lines $\omega=Uk$, showing that the slope of Branch II increases with wind speed but remains systematically smaller than $U$.
In order to quantify $U_{c}$, we subsequently extract the local maximum of $|\hat{a}_{z}(f,k)|$ over wavenumber $k_{c}$ (green circles) for a given frequency, and perform a linear fit along Branch II by $\omega=U_{c}k_{c}$. The resulting convection velocity as a function of wind speed is shown in Fig. 6(d). We see a clear positive correlation between the wind velocity and the velocity $U_{c}$ characterizing the advection of the turbulent pressure fluctuations along the chain. On average, the ratio $U_{c}/U=0.75\pm 0.04$, consistent with 0.8 in our previous study on the reduced chain of pendulums [zhang2024large], and comparable to the convection velocity of wall-pressure fluctuations in a zero-pressure-gradient turbulent boundary layer [choi1990space]. We notice that the ratio $U_{c}/U$ approaches or exceeds 1 for higher wind speeds ($U=6.58$ m/s and 7 m/s), possibly due to the larger scatter of the local maxima in this regime. To summarize, by exploiting the embedded accelerometers of the smartphone chain, we have measured the spatiotemporal Fourier spectrum of the pendulum oscillations and revealed the emergence of a two-branch dispersion relation under wind forcing. Branch I arises from the resonance of individual smartphones, each responding locally to turbulent pressure fluctuations near their natural frequency $\omega_{0}$. Branch II, on the other hand, originates from the collective motion of the chain, where the pendulums respond coherently to the advection of large-scale turbulent structures. These underlying mechanisms explain the coexistence of a nearly flat resonant branch and a linear convective branch in the spectrum. From the latter, we reliably extracted the convection velocity of large-scale turbulent pressure fluctuations. The measured convection velocity scales as a constant fraction of the free-stream velocity, consistent with observations in turbulent boundary layers. This validates our approach as a robust method for large scale measurements of flow-induced oscillations of structures, at dimensions directly comparable to Ned Kahn’s Kinetic Facade artworks.
3.2 Wave buoys